A 7 g metal sample absorbed 211 J of heat as it was heated from 28° C to 85° C. What is the specific heat of the metal?


1
Expert's answer
2021-03-11T08:42:48-0500

Q = 211 J

m = 7g

"\Delta T" = 85⁰C - 28⁰C = 57⁰C

c = specific heat

Q = mc"\Delta T"

211 = 7×c×57

c = 211/(57×7)

c = 0.529 j/g ⁰C